a>b>c, a+b+c=de, de+d+e=fg, fg+f+g=hi. use 1 to 9 only
Answer by greenestamps(13200)



a+b+c = de

The sum of three 1-digit integers is a 2-digit integer. The tens digit of that 2-digit integer MIGHT be 2; but it is much more likely to be 1. So let's assume d=1.

de+d+e = fg

If d=1, then this becomes 1e+1+e = fg, or 11+2*e = fg. That means f has to be 2. So now we know 11+2*e = 2g.

fg+f+g = hi

Since f=2, this becomes 2g+2+g = hi, or 22+2*g = hi. That means h has to be 3. So now we know 22+2*g = 3i.

So digits 1, 2, and 3 are d, f, and h; so a, b, c, e, g, and i must be, in some order, digits 4 through 9.

Now we now 11+2*e = 2g; and g can't be 1, 2, or 3. That means e has to be 7, 8, or 9.

And we know 22+2*g = 3i; and i can't be 1, 2, or 3. That means g has to be 6, 7, 8, or 9.

That's all the logical analysis I can see that we can do. Now we just need to try sets of digits for a, b, and c for which the sum a+b+c is greater than 13 and less than 20; and see where that leads us. For each set of digits we choose for a, b, and c, we stop looking at that case when a letter needs to be a digit that has already been used.
   a, b, c    e (units digit  g (units digit  i (units digit
                 of a+b+c)       of 11+2*e)      of 22+2*g)
  -----------------------------------------------------------
    456              5 X
    457              6               3 X
    458              7               5 X
    459              8               7               6  << YES; this should be the solution; but let's keep checking
    467              7 X
    468              8 X
    469              9 X
    478              9               9 X
    567              8               7 X
    568              9               9 X

There is indeed a single solution: (a,b,c,d,e,f,g,h,i) = (9,5,4,1,8,2,7,3,6)

a>b>c? 9>5>4? YES
a+b+c = de? 9+5+4 = 18? YES
de+d+e=fg? 18+1+8 = 27? YES
fg+f+g=hi? 27+2+7 = 36? YES